# AsynEVO: Asynchronous Event-Driven Visual Odometry for Pure Event Streams

### v2

@@ I Introduction
The ego-motion estimation plays a critical role in the automated robots. Typically, the frame-based camera is adopted to infer the ego-motion from intensity images, which is called Visual Odometry (VO) [1, 2, 3] or Simultaneous Localization and Mapping (SLAM) [4, 5]. Over the past decades, an implicit paradigm has formed where a visual frontend tracks the sparse visual features to establish feature associations, and a parallel backend triangulates landmarks from the related features and solves the Maximum A Posteriori (MAP) problem to infer the ego-motion [1, 5]. However, the traditional frame-based feature tracking and discrete-time MAP estimation methods have a limited temporal resolution for the fixed sampling frequency. Moreover, the frame-based camera captures the instantaneous brightness of the current field view at a fixed frequency, and thus the inter-frame motion information is almost abandoned. Although a camera with high frame rates can attenuate the drawback, the resulting computational consumption due to booming images would be quickly intractable. Therefore, the traditional estimation paradigm for frame-based cameras naturally has a limited temporal resolution.
@@
The rest of this article is organized as follows. Sec. II reviews the relevant literature. In Sec. III, we first overview the whole system framework, and then introduce the design and model of components, respectively. Sec. IV demonstrates the superiority of our proposal on public and own-collected datasets. Finally, the conclusions are drawn in Sec. V.
@@ II Related Work
Researchers have proposed different schemes to realize ego-motion estimation for event cameras, mainly including frame-like, learning-based, and continuous-time methods. Frame-like methods accumulated event streams to create some composite feature structures, such as Surface of Active Events (SAE) and event frame[16, 12]. Sparse features were then detected and tracked in a frame-by-frame manner using enhanced conventional feature-based methods [17, 18]. Subsequently, the discrete-time backend optimization or filtering techniques were applied to estimate the motion states. Some studies presented approaches to enhance the frontend performance by directly combining the inertial measurements [19] or other sensors [20]. As the depth of environments can not be measured by monocular camera, event-based stereo visual odometry was also studied to restore the metric scale [21, 22]. However, frame-like methods primarily treat the event-based estimation problem as a synchronous discrete-time paradigm, which limited their ability to harness the asynchronicity and high-temporal resolution of event cameras.
Learning-based methods, including Convolutional Neural Network (CNN) and Spiking Neural Network (SNN), have been designed to predict optical flow from event streams [23, 24, 25]. A motion estimation module is attached behind the optical flow prediction module to infer the motion states. These CNN-based methods usually process event streams in a synchronous batched manner. Consequently, the asynchronicity of event cameras largely neglected. In contrast, the SNN has an effective mechanism for encoding temporal associations from asynchronous event streams. Some unsupervised learning methods have been suggested to predict both optical flow and ego-motion [24, 25]. Typically, these SNN models require specialized neuromorphic chips for deployment and have inherent difficulty in model training. Moreover, the generalization performance of learning-based methods needs to be improved in future research [23].
CT estimation methods [26, 27, 28, 29] were designed to infer the motion trajectory from asynchronous measurements and unsynchronized sensors, such as those taken from a moving car equipped with a scanning radar [28] or multi-sensor extrinsic calibration [30]. Recently, CT estimation methods have also been applied to event-based trajectory estimation tasks, as the event camera is inherently an asynchronous sensor. For example, Mueggler et al. [11] employed the B-spline to represent the CT trajectory. The estimated motion trajectory is deformed by optimizing the sparse control points to minimize reprojection errors. Subsequently, the inertial measurements were fused to further improve the accuracy and recover the metric scale [13].
@@ III-A System Overview
Figure 1: System pipeline. The first thread inputs event measurements, tracks with asynchronous tracking method, and outputs feature trajectories to a queue. The second thread pops feature trajectories, tries triangulating them, and adds them into the backend graph.
The overall system pipeline primarily consists of two parallel components, i.e., 1) asynchronous event-driven visual frontend, and 2) dynamic sliding-window backend.
In the frontend, the corner features are detected on the SAE. Apart from conventional frame-based methods, the detecting is triggered by the incoming events and is thus asynchronous. When a new feature is detected, we create a data structure, known as feature trajectory, to record the pixel movement of this feature. We attach a separate event-driven tracker to each feature trajectory. The subsequent events, located within the neighbor region, will be tracked in an event-by-event manner. Those feature trajectories that meet preset conditions are fed to the backend.
In the backend, several motion states with the same time interval are initialized to form a sliding-window at system startup. Then, we attempt to triangulate asynchronous feature trajectories within the sliding-window. The corresponding state of each measurement on the feature trajectory is queried with the GP-based interpolation. A landmark will be assigned to the feature trajectory if it is successfully triangulated. We further refine all landmarks and states with a full optimization. After initialization, the incoming feature trajectories are triangulated and add to sliding-window sequentially. If the timestamp of the incoming measurement is larger than the counterpart of the latest state, we execute the dynamic marginalization and the sliding-window optimization. Within the sliding-window, the optimization is defined as a finite factor graph where the adjacent motion states are constrained by the GP prior factor and the event measurement is represented as the GP projection factor.
@@ III-B Asynchronous Visual Frontend
Since the registration table has the same width and height as the resolution of event cameras, it is high-efficient to access and modify elements in the registration table. The search of existing features only pays attention to a small patch within the registration table, and thus has a low time complexity. The small patch also reject the feature clustering and realize sparse features. Similarly, the detection only occurs at the location of the current event measurement and is thus efficient.
As the tracking of active features is triggered by subsequent events, a special case is that a active feature have not be triggered for a long time. For example, a corner feature is detected in noisy events by mistake, and thus it should be discarded promptly. To realize it, we assume the active feature should be triggered frequently within a maximum time threshold. We periodically delete the active features that have not been updated for more than the maximum time threshold. In addition, we also set a minimum time threshold to avoid the feature trajectory collecting a mass of redundant measurements. With patch search and time threshold mechanisms, the result frontend can maintain sparse and asynchronous feature tracking.
@@ III-D Sparse GP priors on $SE(3)$
where $\log(\bullet)$ and $(\bullet)^{\vee}$ are the inverse maps of $\exp(\bullet)$ and $(\bullet)^{\land}$, respectively. Moreover, a sparse covariance matrix [32] for the GP prior residual is given by
$\bm{Q}_{k+1}=\begin{bmatrix}\frac{1}{3}(t_{k+1}-t_{k})^{3}\bm{Q}_{c}&\frac{1}{2}(t_{k+1}-t_{k})^{2}\bm{Q}_{c}\\ \frac{1}{2}(t_{k+1}-t_{k})^{2}\bm{Q}_{c}&(t_{k+1}-t_{k})\bm{Q}_{c}\end{bmatrix}.$ (11)
@@ III-E GP Projection Measurement
where $\bm{\Phi}(t_{\tau},t_{k})$ is the transition matrix [27] and $\bm{Q}_{i,\tau}$ is obtained through substitution of the measurement timestamp $t_{i,\tau}$ into the GP prior covariance matrix (11). Then, we can use (9) to recover $\bm{x}(t_{i,\tau})$. Therefore, the projected pixel position of landmark $\bm{l}_{i}=[x_{i},y_{i},z_{i}]^{\top}$ at timestamp $t_{i,\tau}$ can be written as
@@ 21             $\bm{\chi}_{m}\leftarrow\bm{\chi}_{m}\cup\{\bm{x}_{k}\}$;
The special sparse structure of bundle adjustments plays a pivotal role in maintaining computational efficiency [34]. With our marginalization algorithm, we neither discard any measurements directly nor introduce new constraints between motion states and landmarks. Therefore, the structure of the resulting Hessian matrix retains its arrow-shaped form, implying that the corresponding optimization problem still have a distinct sparse structure and could be solved efficiently. Since the resulting sliding-window can vary with the current data-association condition, we refer to this approach as a dynamic sliding-window or dynamic marginalization.
Based on the factor graph theory, the marginalization is achieved on the linearized Markov blanket of the state variables [35]. Thus, the Markov blanket is first extracted from the original factor graph according to the state variables that should be marginalized, i.e., the outputs of the Alg. 1. Then, the linearization of Markov blanket is achieved by the Schur complement [31]. Finally, the induced marginal distribution on the Markov blanket replaces the corresponding part in the original factor graph as presented in Fig. 3.
III-G Asynchronous System Initialization
For monocular initialization, we first convert the asynchronous initialization to synchronous initialization. Then, the method for conventional discrete systems is leveraged to realize initialization. Concretely, the asynchronous feature trajectories within the sliding-window backend are interpolated at timestamps of motion states. Although the new motion state in our framework can be created with a variable time interval, we choose a fixed time interval in practice for the sake of simplicity. The redundant measurements are discarded directly. After that, the problem becomes a standard discrete monocular initialization. We monitor the average parallax between the first and last motion states. If the parallax condition is satisfied, we adopt five-points method to recover their relative rotation and translation up to a scale. Subsequently, the landmarks are triangulated and the PnP method is utilized to recover the pose of other motion states. Finally, all landmarks and motion states in the sliding-window backend are tuned with a full bundle adjustment.
@@ V Conclusions
In this paper, an asynchronous visual odometry known as AsynVO was proposed to estimate $SE(3)$ trajectory from asynchronous event streams. The proposed method was tested to be credible on both public datasets and real-world experiments.
Although our current version can’t accomplish real-time runtime, the proposed method exhibits potential capacity for future improvements. Since our backend optimization problem becomes partial dense among related poses after the landmark marginalization operation, we can mitigate that by sparsifying the problem through the removal of redundant connections of state variables. Due to the Markov property of the motion state variables, the aforementioned sparsity would not compromise the original optimization problem. Furthermore, our approach can also be readily extended to integrate stereo, inertial, and White-Noise-On-Jerk (WNOJ) configurations in the future, which would be beneficial to inhibiting the scale drift and supporting aggressive trajectories. The additional observations would permit smaller window size and realize real-time estimation.

### v3

@@ I Introduction
The ego-motion estimation plays a critical role in the automated robots [1, 2]. Typically, the frame-based camera is adopted to infer the ego-motion from intensity images, which is called Visual Odometry (VO) [3, 4] or Simultaneous Localization and Mapping (SLAM) [5, 6]. Over the past decades, an implicit paradigm has formed where a visual frontend tracks the sparse visual features to establish feature associations, and a parallel backend triangulates landmarks from the related features and solves the Maximum A Posteriori (MAP) problem to infer the ego-motion [3, 6]. However, the traditional frame-based feature tracking and discrete-time MAP estimation methods have a limited temporal resolution for the fixed sampling frequency. Moreover, the frame-based camera captures the instantaneous brightness of the current field view at a fixed frequency, and thus the inter-frame motion information is almost abandoned. Although a camera with high frame rates can attenuate the drawback, the resulting computational consumption due to booming images would be quickly intractable. Therefore, the traditional estimation paradigm for frame-based cameras naturally has a limited temporal resolution.
@@
The rest of this article is organized as follows. Sec. II reviews the relevant literature. In Sec. III, we first overview the whole system framework, and then introduce the design and model of components, respectively. Sec. IV demonstrates the superiority of our proposal on public and own-collected data. Finally, the conclusions are drawn in Sec. V.
@@ II Related Work
Researchers have proposed different schemes to realize ego-motion estimation for event cameras, mainly including frame-like, learning-based, and continuous-time methods. Frame-like methods accumulated event streams to create some composite feature structures, such as Surface of Active Events (SAE) and event frame[17, 13]. Sparse features were then detected and tracked in a frame-by-frame manner using enhanced conventional feature-based methods [18, 19]. Subsequently, the discrete-time backend optimization or filtering techniques were applied to estimate the motion states. Some studies presented approaches to enhance the frontend performance by directly combining the inertial measurements [20] or other sensors [21]. As the depth of environments can not be measured by monocular camera, event-based stereo visual odometry was also studied to restore the metric scale [22, 23]. However, frame-like methods primarily treat the event-based estimation problem as a synchronous discrete-time paradigm, which limited their ability to harness the asynchronicity and high-temporal resolution of event cameras. Learning-based methods, including Convolutional Neural Network (CNN) and Spiking Neural Network (SNN), have been designed to predict optical flow from event streams [24, 25]. A motion estimation module is attached behind the optical flow prediction module to infer the motion states. These CNN-based methods usually process event streams in a synchronous batched manner. Consequently, the asynchronicity of event cameras largely neglected. In contrast, the SNN has an effective mechanism for encoding temporal associations from asynchronous event streams. Some unsupervised learning methods have been suggested to predict both optical flow and ego-motion [26, 25]. Typically, these SNN models require specialized neuromorphic chips for deployment and have inherent difficulty in model training. Moreover, the generalization performance of learning-based methods needs to be improved in future research [24]. CT estimation methods [27, 28, 29] were designed to infer the motion trajectory from asynchronous measurements and unsynchronized sensors, such as those taken from a moving car equipped with a scanning radar [28] or multi-sensor extrinsic calibration [30]. Recently, CT estimation methods have also been applied to event-based trajectory estimation tasks, as the event camera is inherently an asynchronous sensor. For example, Mueggler et al. [12, 14] employed the B-spline to represent the CT trajectory. The estimated motion trajectory is deformed by optimizing the sparse control points to minimize reprojection errors.
@@ III-A System Overview
The overall system pipeline primarily consists of two parallel components, i.e., 1) asynchronous event-driven visual frontend, and 2) dynamic sliding-window backend. In the frontend, the corner features are detected on the SAE. Apart from conventional frame-based methods, the detecting is triggered by the incoming events and is thus asynchronous. When a new feature is detected, we create a data structure, known as feature trajectory, to record the pixel movement of this feature. We attach a separate event-driven tracker to each feature trajectory. The subsequent events, located within the neighbor region, will be tracked in an event-by-event manner. Those feature trajectories that meet preset conditions are fed to the backend. In the backend, several motion states with the same time interval are initialized to form a sliding-window at system startup. Then, we attempt to triangulate asynchronous feature trajectories within the sliding-window. The corresponding state of each measurement on the feature trajectory is queried with the GP-based interpolation. A landmark will be assigned to the feature trajectory if it is successfully triangulated. We further refine all landmarks and states with a full optimization. After initialization, the incoming feature trajectories are triangulated and add to sliding-window sequentially. If the timestamp of the incoming measurement is larger than the counterpart of the latest state, we execute the dynamic marginalization and the sliding-window optimization. Within the sliding-window, the optimization is defined as a finite factor graph where the adjacent motion states are constrained by the GP prior factor and the event measurement is represented as the GP projection factor.
@@ III-B Asynchronous Visual Frontend
Since the registration table has the same width and height as the resolution of event cameras, it is high-efficient to access and modify elements in the registration table. The search of existing features only pays attention to a small patch within the registration table, and thus has a low time complexity. The small patch also reject the feature clustering and realize sparse features. Similarly, the detection only occurs at the location of the current event measurement and is thus efficient. As the tracking of active features is triggered by subsequent events, a special case is that a active feature have not be triggered for a long time. For example, a corner feature is detected in noisy events by mistake, and thus it should be discarded promptly. To realize it, we assume the active feature should be triggered frequently within a maximum time threshold. We periodically delete the active features that have not been updated for more than the maximum time threshold. In addition, we also set a minimum time threshold to avoid the feature trajectory collecting a mass of redundant measurements. With patch search and time threshold mechanisms, the result frontend can maintain sparse and asynchronous feature tracking.
Figure 1: System pipeline. The first thread inputs event measurements, tracks with asynchronous tracking method, and outputs feature trajectories to a queue. The second thread receives feature trajectories, tries triangulating them, and adds them into the backend graph.
@@ III-D Sparse GP priors on $SE(3)$
where $\log(\bullet)$ and $(\bullet)^{\vee}$ are the inverse maps of $\exp(\bullet)$ and $(\bullet)^{\land}$, respectively.
@@ III-E GP Projection Measurement
where $\bm{\Phi}(t_{\tau},t_{k})$ is the transition matrix [27] and $\bm{Q}_{i,\tau}$ is obtained through substitution of the measurement timestamp $t_{i,\tau}$ into the GP prior covariance matrix (LABEL:Qp). Then, we can use (9) to recover $\bm{x}(t_{i,\tau})$. Therefore, the projected pixel position of landmark $\bm{l}_{i}=[x_{i},y_{i},z_{i}]^{\top}$ at timestamp $t_{i,\tau}$ can be written as
@@ 21             $\bm{\chi}_{m}\leftarrow\bm{\chi}_{m}\cup\{\bm{x}_{k}\}$;
The special sparse structure of bundle adjustments plays a pivotal role in maintaining computational efficiency [33]. With our marginalization algorithm, we neither discard any measurements directly nor introduce new constraints between motion states and landmarks. Therefore, the structure of the resulting Hessian matrix retains its arrow-shaped form, implying that the corresponding optimization problem still have a distinct sparse structure and could be solved efficiently. Since the resulting sliding-window can vary with the current data-association condition, we refer to this approach as a dynamic sliding-window or dynamic marginalization. Based on the factor graph theory, the marginalization is achieved on the linearized Markov blanket of the state variables [34]. Thus, the Markov blanket is first extracted from the original factor graph according to the state variables that should be marginalized, i.e., the outputs of the Alg. 1. Then, the linearization of Markov blanket is achieved by the Schur complement [31]. Finally, the induced marginal distribution on the Markov blanket replaces the corresponding part in the original factor graph as presented in Fig. 3.
@@ V Conclusions
In this paper, an asynchronous visual odometry known as AsynVO was proposed to estimate $SE(3)$ trajectory from asynchronous event streams. The proposed method was tested to be credible on both public datasets and real-world experiments. Although our current version can’t accomplish real-time runtime, the proposed method exhibits potential capacity for future improvements. Since our backend optimization problem becomes partial dense among related poses after the landmark marginalization operation, we can mitigate that by sparsifying the problem through the removal of redundant connections of state variables. Due to the Markov property of the motion state variables, the aforementioned sparsity would not compromise the original optimization problem. Furthermore, our approach can also be readily extended to integrate stereo, inertial, and White-Noise-On-Jerk (WNOJ) configurations in the future, which would be beneficial to inhibiting the scale drift and supporting aggressive trajectories. The additional observations would permit smaller window size and realize real-time estimation.